OpenStudy (anonymous):

is this correct? 7 1/2- 2 7/12= 29/2-31/12 174/12-31/12=143/12

OpenStudy (anonymous):

7 1/2- 2 7/12 =(7x2 +1)/2 -(2x12 +7)/12 =15/2 -31/12 =(15x6 -31)/12 =59/12

OpenStudy (anonymous):

59/12 is the exact?

OpenStudy (anonymous):

what wud the estimte be

OpenStudy (anonymous):

it is the simplest form already

OpenStudy (anonymous):

i know its asking for the estimate also